OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (campbell_st):

you need to find a pair where the 2nd line is a multiple of the 1st line... then the lines will have a infinite number of solutions.

OpenStudy (anonymous):

I don't really get it

OpenStudy (campbell_st):

ok... so if I have 3x - 2y = 12 and -12x + 8y = -48 I multiplied every term in the 1st line by -4 to get the 2nd line. the 1st pair are -4x + 6y = 20 4x - 3y = -2 looking at the 1st line the 4x has been multiplied by -1 to get -4x in the 2nd line applying this rule, then 20y should become -20y in the 2nd line... but its -3y so these 2 lines are different... and will intersect.... hope that makes sense